OpenStudy (anonymous):

A box of cereal A contains 29 grams of carbohydrates per serving. A box of cereal B contains 32 grams of carbohydrates per serving. If two boxes of cereal A are mixed with three boxes of cereal B, what will be the number of grams of carbohydrates per serving?

OpenStudy (anonymous):

S = number of grams of carbohydrates per serving of mixture S = [(2 * 29) + (3 * 32)]/5 S = (58 + 96)/5 S = 154/5 S = 30.8 grams of carbohydrates/serving

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

My pleasure(:

OpenStudy (anonymous):

u are ganna get a medal

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

ur welcome